# 3D Diffusion Policy: Generalizable Visuomotor Policy Learning via Simple 3D Representations

### v2

@@ I Introduction
To comprehensively evaluate DP3, we have developed a simulation benchmark comprising 72 diverse robotic tasks from 7 domains, alongside a real-world benchmark containing 4 challenging manipulation tasks. Our extensive experiments demonstrate that although DP3 is conceptually straightforward, it exhibits several notable advantages over 2D-based diffusion policies and other baselines:
@@
    Safety. An interesting observation in our real-world experiments is that DP3 seldom gives erratic commands in real-world tasks, unlike baseline methods which often do and exhibit unexpected behaviors, posing potential damage to the robot hardware.
@@ II-B Visual Imitation Learning
Imitation learning offers an efficient way for robots to acquire human-like skills, typically relying on extensive observation-action pairs from expert demonstrations. Given the challenges in accurately estimating object states in the real world, visual observations such as images have emerged as a practical alternative. While 2D image-based policies [40, 12, 10, 37, 17, 57, 69, 16] have predominated the field, the significance of 3D is increasingly recognized [61, 82, 80, 15, 14, 30].
@@ II-C Learning Dexterous Skills
Achieving human-like manipulation skills in robots has been a longstanding objective pursued by robotics researchers. Reinforcement learning has been a key tool in this endeavor, enabling robots with dexterous hands to master a variety of tasks, such as pouring water [48, 81], opening doors [50, 22, 8], rotating objects [45, 76, 78, 46], reorienting objects [19, 7, 6], spinning pens [35], grasping tools [1], executing handovers [83, 25], and building Legos [9]. Imitation learning offers another pathway, with approaches like DIME [3] and DexMV [48] translating human hand movements into robotic actions through retargeting and enabling learning from human videos. Our work, however, diverges from these specific design-centric methods. We demonstrate that enabling the acquisition of these complex skills with minimal demonstrations could be achieved by improving the imitation learning algorithm itself.
@@ V-C Generalization
V-D Safety Violation
In our real-world experiments, we surprisingly observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors, which necessitates human termination to ensure robot safety. This situation is defined as safety violation. We compute the safety violation rate in our main real-world experiments, shown in Table XIII. Interestingly, DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning.

### v3

@@ I Introduction
To comprehensively evaluate DP3, we have developed a simulation benchmark comprising 72 diverse robotic tasks from 7 domains, alongside 4 real-world tasks including challenging dexterous manipulation on deformable objects. Our extensive experiments demonstrate that although DP3 is conceptually straightforward, it exhibits several notable advantages over 2D-based diffusion policies and other baselines:
@@
    Safe deployment. An interesting observation in our real-world experiments is that DP3 seldom gives erratic commands in real-world tasks, unlike baseline methods which often do and exhibit unexpected behaviors, posing potential damage to the robot hardware.
@@
In summary, our contributions are four-fold:
  1. 1.
    We propose 3D Diffusion Policy (DP3), an effective visuomotor policy that generalizes across diverse aspects with few demonstrations.
  2. 2.
    To reduce the variance brought by benchmarks and tasks, we evaluate DP3 in a broad range of simulated and real-world tasks, showing the universality of DP3.
  3. 3.
    We conduct comprehensive analyses on visual representations in DP3 and show that a simple point cloud representation might be preferred over other intricate 3D representations.
  4. 4.
    Our real-world experiments demonstrate that complex high-dimensional dexterous manipulation tasks could be handled with simple design and little human data.
@@ II-B Visual Imitation Learning
Imitation learning offers an efficient way for robots to acquire human-like skills, typically relying on extensive observation-action pairs from expert demonstrations. Given the challenges in accurately estimating object states in the real world, visual observations such as images have emerged as a practical alternative. While 2D image-based policies [40, 12, 10, 37, 17, 57, 69, 16] have predominated the field, the significance of 3D is increasingly recognized [61, 83, 81, 15, 14, 30, 70].
@@ II-C Learning Dexterous Skills
Achieving human-like manipulation skills in robots has been a longstanding objective pursued by robotics researchers. Reinforcement learning has been a key tool in this endeavor, enabling robots with dexterous hands to master a variety of tasks, such as pouring water [48, 82], opening doors [50, 22, 8], rotating objects [45, 77, 79, 46], reorienting objects [19, 7, 6], spinning pens [35], grasping tools [1], executing handovers [84, 25], and building Legos [9]. Imitation learning offers another pathway, with approaches like DIME [3], DexMV [48], and DexCap [70] translating human hand movements into robotic actions through retargeting and enabling learning from humans. Our work, however, diverges from these specific design-centric methods. We demonstrate that enabling the acquisition of these complex skills with minimal demonstrations could be achieved by improving the imitation learning algorithm itself.
@@ V-C Generalization
V-D An Interesting Observation: Safe Deployment
In our real-world experiments, we observe that image-based and depth-based diffusion policies often deliver unpredictable behaviors in real-world experiments, which necessitates human termination to ensure robot safety. We define this situation as safety violation and compute the safety violation rate in our main real-world experiments, shown in Table XIII. Interestingly and surprisingly, we find that DP3 rarely violates the safety, showing that DP3 is a practical and hardware-friendly method for real robot learning. An intuitive explanation is that since the robots operate in 3D space, directly observing 3D information helps avoid collision. It is important to note that our assessment of safety is primarily qualitative. We intend to explore a more theoretical understanding of this observation in our future work.